OpenStudy (thats_kyy):

Jamie is 5 years older than Ella. Jamie's age is 11 years less than three times Ella's age. The system below models the relationship between Jamie's age (j) and Ella's age (e): j = e + 5 j = 3e − 11 Which of the following methods is correct to find Jamie's and Ella's age?

OpenStudy (thats_kyy):

- Solve j + 5 =3j to find the value of j - Write the points where the graphs of equations intersect the x-axis - Write the points where the graphs of equations intersect the y-axis - Sole e + 5 = 3e - 11 to find the value of e

OpenStudy (thats_kyy):

@robtobey

OpenStudy (thats_kyy):

@dan815

OpenStudy (thats_kyy):

@perl

OpenStudy (thats_kyy):

@texaschic101

OpenStudy (anonymous):

Solve e + 5 = 3e - 11 to find the value of e

OpenStudy (thats_kyy):

Thats The Answer?

OpenStudy (anonymous):

e = 8 The problem statement does require some plotting.